OpenStudy (anonymous):

rewrite with a sinle exponent (9^7)^3

OpenStudy (angela210793):

=9^(7*3)=9^21

OpenStudy (jamesj):

In general \[(a^x)^y = a^{xy}.\] Hence the result people are giving you above.

OpenStudy (anonymous):

??

OpenStudy (jamesj):

For example \[(2^5)^3 = 2^{5 \ times \ 3} = 2^{15}\] In the example I just did, a = a, x = 5 and y = 3 Beginning to make sense?

OpenStudy (jamesj):

You can see this explicitly if you write it out. What does (2^5)^3 mean? It means first of all 2^5 multiplied out three times \[(2^5)^3 = 2^5 . 2^5 . 2^5\] Now this is equal to \[2^{5 + 5 + 5} = 2^{5 \ \times \ 3} = 2^{15}\]